# Paper/main_online_appendix

### v3

@@ 3.3.1 Lottery Reversals
(Classic preference reversals). Suppose the DM is weakly risk-averse ($u$ is concave), and again consider the simple lotteries $x=(\$23.5,0.19)$, $y=(\$4.75,0.94)$. First, consider binary choice between the (equal expected value) lotteries. Since any risk-averse DM weakly prefers $y$ to $x$, our model predicts that the DM is more likely to choose $y$.
@@
Figure 3(a) plots the expected certainty equivalents $\mathbb{E}[CE(l,Z)]$ simulated from our model for simple lotteries $l$ with the same expected value as $x$ and $y$. Notice that the high-risk lottery $x$ is valued higher than the low-risk lottery $y$ on average, despite the fact that $y$ is weakly preferred to $x$ for our risk-averse DM. Intuitively, $x$ is dissimilar to and therefore difficult to compare to money, so its valuation is inflated toward the midpoint of the undominated range of prices $[0,w_{x}]$. On the other hand, $y$ is easier to compare to money, so its valuation is less distorted, and if anything is pulled downward toward the midpoint of undominated prices $[0,w_{y}]$. We have a reversal: $\rho(y,x)\geq 1/2$ and yet $\mathbb{E}[CE(x,Z)]>\mathbb{E}[CE(y,Z)]$.
@@
(Reversals with probability equivalents). Consider the lotteries $x=(\$23.50,19\%)$ and $y=(\$4.75,94\%)$ from Example 1. Instead of valuing $x$ and $y$ against money, suppose the DM assesses their probability-equivalents: the probability $p$ that makes the lottery $z=(\$24,p)$ indifferent to each. Whereas $y$ was easier to compare to money, $x$ is easier to compare to this new numeraire, which is more similar to $x$ than to $y$. Our model predicts that this change in numeraire reverses the distortions in the valuation of $x$ and $y$ [^17] Slovic et al., (2002) experimentally study a related procedure of “probability matching”, in which subjects indicate the probability $p$ that makes $x^{\prime}=(\$23.5,p)$ indifferent to $y$, from which a preference between $x$ and $y$ can be inferred. This can be seen as a special case of our manipulation in which the price list is made perfectly comparable to the high risk lottery $x$, i.e. $z=(\$23.50,p)$. They document that the inconsistency with direct choice disappears when high-risk/low-risk lottery pairs are assessed using probability matching, consistent with the predictions of our model.
@@ 3.3.2 Intertemporal Reversals
[…]
(a) Present Value Equivalents $\mathbb{E}[PVE(l,z)]$
(b) Time Equivalents $\mathbb{E}[T(l,z)]$
@@
As we saw in lotteries, the direction of these distortions may be reversed by changing the units of valuation. Suppose the DM values options in terms of time-equivalents: the time $t$ that makes the delayed payment (\$$(\$27.50,t)$ indifferent to $x$ or $y$. Formally, the DM values $\upsilon$ against a time list $Z=\{z^{1},...,z^{n}\}$, where each $z^{k}=(\$27.50,t_{k})$. As $x$ is more similar to the numeraire under this valuation mode, the model predicts the distortions in time equivalents will be flipped relative to present value equivalents, as Figure 4(b) illustrates. Specifically, we have $\mathbb{E}[T(y,Z)]<\mathbb{E}[T(x,Z)]$ and so the “reversal” relative to direct choice disappears.
@@ 3.4 Biases in Valuation of Risk and Time
3.4.1 Probability-Weighting in Lottery Valuation
@@
Consider the standard paradigm used to estimate the probability weighting function, in which the DM provides certainty equivalents of simple lotteries $l=(\overline{w},p_{l})$. Note that for $p_{l}$ sufficiently close to 0 or 1, $l$ is easy to compare to prices and so valuations are close to accurate, whereas for interior values of $p_{l}$, tradeoffs between payoff amounts and probabilities make $l$ harder to compare to prices, producing pull-to-center distortions: small probabilities are overvalued and large probabilities are undervalued. These two facts generate apparent inverse-S probability weighting in the DM’s valuations, as seen in Figure 5(a), which plots our model’s predicted normalized certainty equivalents $\mathbb{E}[CE(l,Z)]/\overline{w}$ as a function of $p_{l}$.
@@
Importantly, our model does not predict that probability weighting occurs generically in choice; instead, it results from the difficulty of comparing a lottery to a price list of certain payments. This is important for two reasons. First, our model predicts probability weighting in valuation even absent such distortions in binary choice, which is consistent with evidence that the inverse S-shaped probability weighting function is far more prominent in valuation tasks than in direct choice (Harbaugh et al.,, 2010; Bouchouicha et al.,, 2023). Second, we predict that these patterns of probability weighting will be sensitive to the specific valuation paradigm. In particular, we predict that it is possible to reverse the pattern of apparent probability-weighting with an appropriate choice of price list currency.
@@
Consider an alternative paradigm for estimating the probability weighting function, in which the DM provides probability equivalents of a certain payment: the probability $p$ that makes the lottery $z=(\overline{w},p)$ indifferent to $c=(w_{c},1)$. Tracing out the probability equivalents $p$ as a function of the normalized certain payments $w_{c}/\overline{w}$ should — absent complexity-driven distortions — reveal the same preferences as the certainty equivalents discussed above. Figure 5(b) plots the predicted relationship between the certain payment amount $w_{c}/\overline{w}$ (y-axis) and the associated probability equivalent $\mathbb{E}[PE(c,Z)]$ (x-axis), in which the certain payment $c=(w_{c},1)$ is valued against a probability list $Z=\{z^{1},...,z^{n}\}$, for $z^{k}=(\overline{w},p_{k})$. Here we see a reversal of the inverse S-shaped pattern: the difficulty of comparing sure payments against the numeraire good causes probability equivalents to be compressed towards the middle of the price list, generating apparent underweighting of small probabilities and overweighting of large probabilities. [^19] While Sprenger, (2015) shows that a model of stochastic reference points can predict a larger degree of risk aversion in probability vs. certainty equivalents in this setting, this account cannot explain the simultaneous apparent overweighting of large probabilities that our model predicts in probability equivalents.
3.4.2 Hyperbolic Discounting in Intertemporal Valuation
@@
Now, consider a standard paradigm used to estimate time discounting: the DM values delayed payments $\upsilon=(\overline{m},t_{\upsilon})$ in terms of money today. Following the same logic as in lottery valuation, pull-to-center effects result in payments closer to the present to be undervalued and payments further in the future to be overvalued, with valuations becoming accurate as $t_{\upsilon}\to 0$. This produces complexity-driven hyperbolic discounting, consistent with evidence from Enke et al., (2023). We further predict that this pattern persists when front-end delays are incorporated, since the hyperbolicity predicted by our model is not driven by a preference for the present. This is also supported by evidence from Enke et al., (2023).
@@
Figure 6(a) plots the normalized valuations $\mathbb{E}[PVE(\upsilon,Z)]/\overline{m}$ as a function of the delay $t_{\upsilon}$, where $\upsilon$ is valued against a price list $Z=\{z^{1},...,z^{n}\}$ of immediate payments adapted to $\upsilon$. As the figure shows, the DM’s valuations display apparent hyperbolicity: she undervalues payments close to the present and overvalues payments with longer delays.
@@
The pull-to-center distortions in our model echo the predictions of Bayesian cognitive noise models, in which cognitive noise compresses the DM’s action toward an intermediate default. Unlike our model, these accounts cannot explain why these compression effects disappear near boundary parameters (i.e. $p_{l}=0,1$ in lottery valuation, and $t_{\upsilon}=0$ in intertemporal valuation) without the auxiliary assumption that cognitive noise vanishes at these boundaries (Enke and Graeber,, 2023; Enke et al.,, 2023). As our model explains why options are easier to compare to prices at these boundaries, it can be thought of as providing theoretical justification for such an assumption.
@@ 4.1 Tests of Complexity Measures
The purpose of these experiments is two-fold. First, we show that our proposed complexity measures indeed capture the difficulty of comparisons. We consider three natural indicators of choice complexity: choice errors, choice inconsistency, and subjective uncertainty, and test whether they are decreasing in the value-dissimilarity ratio, as our theory predicts. Second, we show that our model captures quantitatively important features of choice not explained by standard models. In each domain, we structurally estimate our model and compare its performance against leading behavioral models.
We carry out these analyses in three parallel binary choice datasets. We run new experiments in multiattribute and intertemporal choice and compile existing data from Enke and Shubatt, (2023) and Peterson et al., (2021) to study lottery choice. In our experiments, we recruit participants through an online survey platform to make 50 incentivized binary choice problems. For each problem, we elicit participants’ subjective certainty in the optimality of their decision. In order to measure choice consistency, 10 of these problems are randomly repeated throughout the survey. We collect an average of 37 choices for each of 662 multiattribute choice problems and 1,097 intertemporal problems – a total of more than 66,000 individual decisions. The compiled risk dataset includes nearly 10,000 binary choice problems (over 1 million decisions) and includes similar measures of cognitive uncertainty and choice consistency. [^20] Cognitive uncertainty is elicited only for the 500 problems from Enke and Shubatt, (2023). Problems are only repeated in the Peterson et al., (2021) experiment. Unlike our experiments, subjects see these repeated problems immediately after giving an initial response.
@@
Figure 7: Binscatter of problem-level error rates, choice inconsistency, and cognitive uncertainty versus the value-dissimilarity ratio. Problem-level error rates in panels (a), (d), and (g) are constructed as the rate of choosing (a) the lower-value option, (d) the less-preferred option according to the best-fit exponential discounted utility model, and (g) the less-preferred option according to the best fit expected utility model (see Appendix E for estimation details). Problem-level inconsistency rates in panels (b), (e), and (h) are constructed as the percent chance subjects choose differently in any repeated instance of the choice problem. Problem-level cognitive uncertainty in panels (c), (f), (i) is constructed as the average subjective likelihood that subjects assign to making an error in that choice problem, i.e. choosing the less-preferred option. The value-dissimilarity ratios $R_{L1}$, $R_{CPF}$, and $R_{CDF}$ follow Definitions 1, 4, and 3, respectively, where the preference parameters $\beta$, $\delta$, and $u$ are given by the known attribute weights in multiattribute choice, the best-fit exponential discount rate in intertemporal choice, and the best-fit expected utility preferences in lottery choice, respectively.
@@
Figures 7(a), 7(b), and 7(c) show binned scatterplots relating the $L_{1}$ ratio (x-axis) to choice errors, cognitive uncertainty, and inconsistency. All three complexity indicators are strongly decreasing in the ratio. The average error rate, around 5% for problems near-dominance, increases five-fold for problems with the lowest values of the $L_{1}$ ratio ($R^{2}=0.32$). We take this as strong evidence that (1) the ratio indeed captures choice complexity, and (2) decision-makers respond to this complexity by making more random choices. Importantly, these relationships are not merely driven by variation in value differences: they are unchanged when controlling for value differences, as the regression analysis in Appendix Table 3 shows. [^21] We pre-registered analyses restricting to subjects who do not report using a calculator in the experiment (82.5% of the sample), reported in Appendix Table 4. Quantitative relationships are virtually unchanged.
@@
Unlike in multiattribute choice, we cannot directly observe choice errors: the definition of an “error” depends on the decision-maker’s discount function. The CPF ratio also depends on this discount function, so we proceed by estimating a representative-agent exponential discount factor $\delta$ from the choice data (monthly $\hat{\delta}=0.94$). [^22] We also run analyses in which errors are classified using individually estimated discount factors $\hat{\delta}_{i}$ with similar results. See Appendix Figure 14 and Table 7. For details on the specification, see Appendix E.2. Importantly, the discounting estimated from our choice data, which involves choices over money, should not be interpreted as subjects’ “pure” time preferences (Cohen et al.,, 2020): it could reflect, for instance, access to outside credit or beliefs about repayment risk. Rather than identifying time preferences, we are interested in studying which problems are hard given subjects’ required rate of return in our experimental setting.
@@
Figures 7(d), 7(e), and 7(f) relate the CPF ratio (x-axis) to our indicators of choice complexity, coding a choice as an “error” if the subject chooses the option with lower value according to $\hat{\delta}$. Again, we find strong relationships between the ratio and choice “errors,” cognitive uncertainty, and inconsistency. The average “error” rate ranges from around 5% for problems with the highest value of the CPF ratio to 50% for problems with the lowest value of the CPF ratio ($R^{2}=0.6)$. As in multiattribute choice, these relationships are unchanged when controlling for the value difference (see Appendix Table 6). This further indicates that the value-dissimilarity ratio, rather than the value difference alone, drives these relationships.
@@
Lottery Choice. In both lottery choice experiments, participants are asked to choose between two lotteries which pay off different amounts with known probabilities. If selected to receive a bonus, they receive the outcome of a lottery they chose in a randomly selected decision. As in intertemporal choice, both the CDF ratio and our notion of choice “errors” depend on an unknown preference object – the Bernoulli utility function. We proceed by estimating a representative-agent CRRA utility function with additive logit noise (see Appendix E.3 for details), and code “errors” as departures from this estimated model. [^23] We also run all analyses in which errors are classified using individually estimated risk aversion parameters. We restrict this analysis to the Enke and Shubatt, (2023) data since the data in Peterson et al., (2021) contains a low amount of unique choice problems for each subject (75% of subjects face $\leq 14$ unique problems). Results are similar – see Appendix Figure 15 and Table 10. Figures 7(g), 7(h), and 7(i) show the results. Once again, all three outcomes are strongly decreasing in the ratio; in particular, the CDF ratio achieves an $R^{2}$ of 0.45 in variance explained over error rates. Consistent with our results in our other two domains, these relationships are driven by the value-dissimilarity ratio, as opposed to the value difference alone (see Appendix Table 9).

### v4

@@ 3.3.1 Lottery Reversals
(Classic preference reversals). Suppose the DM is weakly risk-averse ($u$ is concave), and again consider the simple lotteries x=(\$23.5,0.19)x=(\\$23.5,0.19), y=(\$4.75,0.94)y=(\\$4.75,0.94). First, consider binary choice between the (equal expected value) lotteries. Since any risk-averse DM weakly prefers $y$ to $x$, our model predicts that the DM is more likely to choose $y$.
@@
Figure 2(a) plots the expected certainty equivalents $\mathbb{E}[CE(l,Z)]$ simulated from our model for simple lotteries $l$ with the same expected value as $x$ and $y$. Notice that the high-risk lottery $x$ is valued higher than the low-risk lottery $y$ on average, despite the fact that $y$ is weakly preferred to $x$ for our risk-averse DM. Intuitively, $x$ is dissimilar to and therefore difficult to compare to money, so its valuation is inflated toward the midpoint of the undominated range of prices $0,w_{x}]$. On the other hand, $y$ is easier to compare to money, so its valuation is less distorted, and if anything is pulled downward toward the midpoint of undominated prices $0,w_{y}]$. We have a reversal: $\rho(y,x)\geq 1/2$ and yet $\mathbb{E}[CE(x,Z)]>\mathbb{E}[CE(y,Z)]$.
@@
(Reversals with probability equivalents). Consider the lotteries $x=(\$23.50,19\%)$ and $y=(\$4.75,94\%)$ from Example 1. Instead of valuing $x$ and $y$ against money, suppose the DM assesses their probability-equivalents: the probability $p$ that makes the lottery z=(\$24,p)z=(\\$24,p) indifferent to each. Whereas $y$ was easier to compare to money, $x$ is easier to compare to this new numeraire, which is more similar to $x$ than to $y$. Our model predicts that this change in numeraire reverses the distortions in the valuation of $x$ and yy. [^18] Slovic et al., (2002) experimentally study a related procedure of “probability matching”, in which subjects indicate the probability $p$ that makes $x^{\prime}=(\$23.5,p)$ indifferent to $y$, from which a preference between $x$ and $y$ can be inferred. This can be seen as a special case of our manipulation in which the price list is made perfectly comparable to the high risk lottery $x$, i.e. z=(\$23.50,p)z=(\\$23.50,p). They document that the inconsistency with direct choice disappears when high-risk/low-risk lottery pairs are assessed using probability matching, consistent with the predictions of our model.
@@ 3.3.2 Intertemporal Reversals
Formally, the DM faces a valuation task $(\upsilon,Z)$, where $\upsilon=(m_{\upsilon},t_{\upsilon})$ is a delayed payment that pays out $m_{\upsilon}>0$ at time $t_{\upsilon}>0$, valued against a price list $Z=\{z^{1},...,z^{n}\}$, where each $z^{k}=(m_{k},0)$ is an immediate payment. This setting is similar to those described in Examples 1 and 2, so we reserve further details for Appendix B.6. Figure 3(a) plots the present value equivalents $\mathbb{E}[PVE(\upsilon,Z)]$ simulated from our model for delayed payments $\upsilon$ with the same present value as $x$ and $y$ (assuming a monthly discount factor $\delta=0.95$). Here, the high-delay option $x$ has a higher valuation than $y$, despite the fact that $\rho(y,x)\geq 1/2$. [^19] Using choice vignettes, Tversky et al., (1990) document similar intertemporal preference reversals.
(a) Present Value Equivalents $\mathbb{E}[PVE(\upsilon,Z)]$
(b) Time Equivalents $\mathbb{E}[T(\upsilon,Z)]$
@@
As we saw in lotteries, the direction of these distortions may be reversed by changing the units of valuation. Suppose the DM values options in terms of time-equivalents: the time $t$ that makes the delayed payment (\$27.50,t)(\\$27.50,t) indifferent to $x$ or $y$. Formally, the DM values $\upsilon$ against a time list $Z=\{z^{1},...,z^{n}\}$, where each $z^{k}=(\$27.50,t_{k})$. As $x$ is more similar to the numeraire under this valuation mode, the model predicts the distortions in time equivalents will be flipped relative to present value equivalents, as Figure 3(b) illustrates. Specifically, we have $\mathbb{E}[T(y,Z)]<\mathbb{E}[T(x,Z)]$ and so the “reversal” relative to direct choice disappears.
@@ 3.4 Biases in Valuation of Risk and Time
3.4.1 Apparent Probability Weighting in Lottery Valuation
@@
Consider the standard paradigm used to estimate risk preferences, in which the DM provides certainty equivalents of simple lotteries $l=(\overline{w},p_{l})$. Note that for $p_{l}$ sufficiently close to 0 or 1, $l$ is easy to compare to prices and so valuations are close to accurate, whereas for interior values of $p_{l}$, tradeoffs between payoff amounts and probabilities make $l$ harder to compare to prices, producing pull-to-center distortions: small probabilities are overvalued and large probabilities are undervalued. These two forces generate apparent inverse-S probability weighting in the DM’s valuations, as seen in Figure 4(a), which plots our model’s predicted normalized certainty equivalents $\mathbb{E}[CE(l,Z)]/\overline{w}$ as a function of $p_{l}$.
@@
Importantly, our model does not predict that probability weighting occurs generically in choice; instead, it results from the difficulty of comparing a lottery to a price list of certain payments. This is important for two reasons. First, our model can generate probability weighting in valuation even absent such distortions in binary choice, which is consistent with evidence that the fourfold pattern of probability weighting is far more prominent in valuation tasks than in direct choice (Harbaugh et al.,, 2010; Bouchouicha et al.,, 2023). Second, we predict that these patterns of probability weighting will be sensitive to the specific valuation paradigm. In particular, we predict that it is possible to reverse the pattern of apparent probability-weighting with an appropriate choice of price list currency.
@@
Consider an alternative paradigm for estimating risk preferences, in which the DM provides probability equivalents of a certain payment: the probability $p$ that makes the lottery $z=(\overline{w},p)$ indifferent to $c=(w_{c},1)$. Tracing out the probability equivalents $p$ as a function of the normalized certain payments $w_{c}/\overline{w}$ should — absent complexity-driven distortions — reveal the same preferences as the certainty equivalents discussed above. Figure 4(b) plots the predicted relationship between the certain payment amount $w_{c}/\overline{w}$ (y-axis) and the associated probability equivalent $\mathbb{E}[PE(c,Z)]$ (x-axis), in which the certain payment $c=(w_{c},1)$ is valued against a probability list $Z=\{z^{1},...,z^{n}\}$, for $z^{k}=(\overline{w},p_{k})$. Here we see a reversal of the inverse S-shaped pattern: the difficulty of comparing sure payments against the numeraire good causes probability equivalents to be compressed towards the middle of the price list, generating apparent underweighting of small probabilities and overweighting of large probabilities. [^20] While Sprenger, (2015) shows that a model of stochastic reference points can predict a larger degree of risk aversion in probability vs. certainty equivalents, this account cannot explain the overweighting of large probabilities that our model predicts in probability vs. certainty equivalents. Feldman and Ferraro, (2024) shows that by allowing utility curvature to differ across valuation procedures, a model of stochastic reference points can generate a difference between probability vs. certainty equivalents similar to our model predictions.
3.4.2 Apparent Hyperbolic Discounting in Intertemporal Valuation
@@
Consider a standard paradigm used to estimate discounting over money: the DM values delayed payments $\upsilon=(\overline{m},t_{\upsilon})$ in terms of money today. Following identical logic as in lottery valuation, pull-to-center effects result in payments closer to the present to be undervalued and payments further in the future to be overvalued, with valuations becoming accurate as $t_{\upsilon}\to 0$. This produces complexity-driven hyperbolic discounting (Enke et al.,, 2025).
@@
To illustrate, Figure 5(a) plots the normalized valuations $\mathbb{E}[PVE(\upsilon,Z)]/\overline{m}$ as a function of the delay $t_{\upsilon}$, where $\upsilon$ is valued against a price list $Z=\{z^{1},...,z^{n}\}$ of immediate payments adapted to $\upsilon$. The DM’s valuations exhibit apparent hyperbolicity: she undervalues payments close to the present and overvalues payments with longer delays. As this pattern is driven by tradeoff complexity rather than a preference for the present, our model also predicts that hyperbolicity persists when front-end delays are incorporated, and so reconciles a puzzle in the literature wherein valuations of delayed payments exhibit marked hyperbolicity and yet are stationary with respect to front-end delays (Cohen et al.,, 2020). [^21] See Supplemental Appendix J for details.
@@
The pull-to-center distortions in our model echo the predictions of Bayesian cognitive noise models, in which noise compresses the DM’s action toward an intermediate default. Unlike our model, these accounts cannot explain why these compression effects disappear near boundary parameters (i.e. $p_{l}=0,1$ in lottery valuation, and $t_{\upsilon}=0$ in intertemporal valuation) without the auxiliary assumption that cognitive noise vanishes at these boundaries (Enke and Graeber,, 2023; Enke et al.,, 2025). As our model explains why options are easier to value at these boundaries, it provides a theoretical basis for such an assumption.
@@ 4.1 Tests of Complexity Measures
The purpose of these experiments is two-fold. First, we show that our proposed complexity measures indeed capture the difficulty of comparisons. We consider three natural indicators of choice complexity: choice errors, choice inconsistency, and subjective uncertainty, and show that they are decreasing in the value-dissimilarity ratio, as our theory predicts. Second, we show that our model captures quantitatively important features of choice not explained by standard models. In each domain, we structurally estimate our model and compare its performance against leading behavioral models.
We carry out these analyses in three parallel binary choice datasets. We run new experiments in multiattribute and intertemporal choice and compile existing data from Enke and Shubatt, (2023) and Peterson et al., (2021) to study lottery choice. In our experiments, we recruit participants through an online survey platform to make 50 incentivized binary choices. For each problem, we elicit participants’ subjective certainty in the optimality of their decision. In order to measure choice consistency, 10 of these problems are randomly repeated throughout the survey. We collect an average of 37 choices for each of 662 multiattribute choice problems and 1,097 intertemporal problems – a total of more than 66,000 individual decisions. The compiled risk dataset includes nearly 10,000 binary choice problems (over 1 million decisions) and includes similar measures of cognitive uncertainty and choice consistency. [^22] Cognitive uncertainty is elicited only for the 500 problems from Enke and Shubatt, (2023). Problems are only repeated in the Peterson et al., (2021) experiment. Unlike our experiments, subjects see these repeated problems immediately after giving an initial response.
@@
Figure 6: Binscatter of problem-level error rates, choice inconsistency, and cognitive uncertainty versus the value-dissimilarity ratio. Problem-level error rates in panels (a), (d), and (g) are constructed as the rate of choosing (a) the lower-value option, (d) the less-preferred option according to the best-fit exponential discounted utility model, and (g) the less-preferred option according to the best fit expected utility model (see Appendix E for estimation details). Problem-level inconsistency rates in panels (b), (e), and (h) are constructed as the percent chance subjects choose differently in any repeated instance of the choice problem. Problem-level cognitive uncertainty in panels (c), (f), (i) is constructed as the average subjective likelihood that subjects assign to making an error in that choice problem, i.e. choosing the less-preferred option. The value-dissimilarity ratios $R_{L1}$, $R_{CPF}$, and $R_{CDF}$ follow Definitions 1, 4, and 3, respectively, where the preference parameters $\beta$, $\delta$, and $u$ are given by the known attribute weights in multiattribute choice, the best-fit exponential discount rate in intertemporal choice, and the best-fit expected utility preferences in lottery choice, respectively. See Appendix Tables 3, 6, and 9 for corresponding regression analyses.
@@
Figures 6(a), 6(b), and 6(c) show binned scatterplots relating the $L_{1}$ ratio (x-axis) to choice errors, cognitive uncertainty, and inconsistency. All three complexity indicators are strongly decreasing in the ratio. The average error rate, around 5% for problems near-dominance, increases five-fold for problems with the lowest values of the $L_{1}$ ratio ($R^{2}=0.32$). We take this as strong evidence that (1) the ratio indeed captures choice complexity, and (2) decision-makers respond to this complexity by making more random choices. Importantly, these relationships are not merely driven by variation in value differences: they are unchanged when controlling for value differences, as the regression analysis in Appendix Table 3 shows. [^23] We pre-registered analyses restricting to subjects who do not report using a calculator in the experiment (82.5% of the sample), reported in Appendix Table 4. Quantitative relationships are virtually unchanged.
@@
Unlike in multiattribute choice, we cannot directly observe choice errors: the definition of an “error” depends on the decision-maker’s discount function. The CPF ratio also depends on this discount function, so we proceed by estimating a representative-agent exponential discount factor $\delta$ from the choice data (monthly $\hat{\delta}=0.94$). For details on the specification, see Appendix E.2. Importantly, the discounting estimated from our choice data, which involves choices over money, should not be interpreted as subjects’ “pure” time preferences (Cohen et al.,, 2020): it could reflect, for instance, access to outside credit or beliefs about repayment risk. Rather than identifying time preferences, we are interested in studying which problems are hard given subjects’ required rate of return in our experimental setting.
@@
Figures 6(d), 6(e), and 6(f) relate the CPF ratio (x-axis) to our indicators of choice complexity, coding a choice as an “error” if the subject chooses the option with lower value according to $\hat{\delta}$. Again, we find strong relationships between the ratio and choice “errors,” cognitive uncertainty, and inconsistency. The average “error” rate ranges from around 5% for problems with the highest value of the CPF ratio to 50% for problems with the lowest value of the CPF ratio ($R^{2}=0.6)$. We also run analyses in which errors are classified using individually estimated discount factors $\hat{\delta}_{i}$ with similar results (see Appendix Figure 12 and Table 7). As in multiattribute choice, these relationships are unchanged when controlling for the value difference (see Appendix Table 6). This further indicates that the value-dissimilarity ratio, rather than the value difference alone, drives these relationships.
@@
Lottery Choice. In both lottery choice experiments, participants are asked to choose between two lotteries which pay off different amounts with known probabilities. If selected to receive a bonus, they receive the outcome of a lottery they chose in a randomly selected decision. As in intertemporal choice, both the CDF ratio and our notion of choice “errors” depend on an unknown preference object – the Bernoulli utility function. We proceed by estimating a representative-agent CRRA utility function with additive logit noise (see Appendix E.3 for details), and code “errors” as departures from this estimated model. Figures 6(g), 6(h), and 6(i) show the results. Once again, all three outcomes are strongly decreasing in the ratio; in particular, the CDF ratio achieves an $R^{2}$ of 0.45 in variance explained over error rates. We additionally run analyses using individually estimated risk aversion parameters to classify choices as errors and find similar results (see Appendix Figure 13 and Table 10). [^24] We restrict this analysis to the Enke and Shubatt, (2023) data since the Peterson et al., (2021) data contain a small number of unique choice problems per subject (75% of subjects face $\leq 14$ unique problems). As in our other domains, these relationships are driven by the value-dissimilarity ratio, as opposed to the value difference alone (see Appendix Table 9).